# TinyGPT-V: Efficient Multimodal Large Language Model via Small Backbones

### v1

@@ Abstract
In the era of advanced multimodel learning, multimodal large language models (MLLMs) such as GPT-4V have made remarkable strides towards bridging language and visual elements. However, the closed-source nature and considerable computational demand present notable challenges for universal usage and modifications. This is where open-source MLLMs like LLaVA and MiniGPT-4 come in, presenting groundbreaking achievements across tasks. Despite these accomplishments, computational efficiency remains an unresolved issue, as these models, like LLaVA-v1.5-13B, require substantial resources. Addressing these issues, we introduce TinyGPT-V, a new-wave model marrying impressive performance with commonplace computational capacity. It stands out by requiring merely a 24G GPU for training and an 8G GPU or CPU for inference. Built upon Phi-2, TinyGPT-V couples an effective language backbone with pre-trained vision modules from BLIP-2 or CLIP. TinyGPT-V’s 2.8B parameters can undergo a unique quantisation process, suitable for local deployment and inference tasks on 8G various devices. Our work fosters further developments for designing cost-effective, efficient, and high-performing MLLMs, expanding their applicability in a broad array of real-world scenarios. Furthermore this paper proposed a new paradigm of Multimodal Large Language Model via small backbones. Our code and training weights are placed at: https://github.com/DLYuanGod/TinyGPT-V and https://huggingface.co/Tyrannosaurus/TinyGPT-V respectively.
@@ 1 Introduction
Recently, with the advent of GPT-4V, an expansive multimodal large language model (MLLM), we’ve seen some impressive capabilities in vision-language understanding and generation yang2023dawn . That being said, it’s pivotal to acknowledge that GPT-4V hasn’t been released open-source, thus restricting universal usage and independent modifications. On the bright side, there has been a recent surge in open-source MLLMs, such as LLaVA and MiniGPT-4, which have demonstrated groundbreaking prowess in some tasks, surpassing GPT-4V in areas like image captioning (IC), visual question answering (VQA), and referring expression comprehension (REC) dai2023instructblip ; liu2023improvedllava ; liu2023llava ; zhu2023minigpt . For example, when put to the test on various visual grounding and question answering tasks, MiniGPT-v2 chen2023minigptv2 emerges as a superior force compared to other conventional vision-language models.
Regardless of the substantial vision language capabilities exhibited by some open-source MLLMs, they still consume an excessive amount of computational resources during both training and inference stages. For example, LLaVA-v1.5-13B liu2023improvedllava used 8 A100 GPUs with 80GB memory over the course of 25.5 hours of training. Because the performance of large language models directly impacts the capabilities of MLLMs, their usage, such as LLaVA-v1.5-13B utilizing Vicuna-13b-v1.5 zheng2023judging , and MiniGPT-v2 leveraging LLaMA2-7B-Chat touvron2023llama , necessitates a substantial number of parameters of large language models to enhance performance in complex tasks like IC, VQA etc zhu2023minigpt . Therefore, we require a large language model that can rival the performance of models such as LLaMA2 and Vicuna-v1.5, without the need for excessive GPU computational resources.
So a new model called TinyGPT-V is proposed which requires just 24G of GPU for training and only 8G of GPU or CPU for inference. It utilizes an advanced large language model, Phi-2 javaheripi2023phi , which is constructed upon Phi li2023textbooks , and reported to have outperformed the best effects of the 13B language models and it demonstrated similar or better results than models that are up to 25 times larger in scale. Regarding visual perception, we make use of the same pre-trained vision modules from BLIP-2 li2023blip2 or CLIP radford2021learning , which incorporates a ViT dosovitskiy2021image as vision encoder as well as a mapping module. Following the training approach of MiniGPT, TinyGPT-V only applies the mapping module between the visual encoder and language model throughout the training process, while freezing all other parameters. TinyGPT-V uses the same dataset as MiniGPT-v2 in various stages of training such as LAION schuhmann2021laion400m , Conceptual Captions changpinyo2021conceptual ; sharma-etal-2018-conceptual , SBU NIPS2011_5dd9db5e etc. lin2015microsoft ; schwenk2022okvqa ; hudson2019gqa ; kiela2020hateful ; lu2021iconqa ; gurari2018vizwiz ; mao2016generation ; kazemzadeh2014referitgame ; yu2016modeling .
In our study, we observed that TinyGPT-V displays many qualities that mirror those present in GPT-4, benefiting greatly from the application of the Phi-2 model. Boasting just 2.8B parameters, TinyGPT-V’s unique quantisation process makes it suitable for local deployment and inference tasks on 8G mobile devices. TinyGPT-V represents a substantial stride in reaching the equilibrium between unparalleled performance and maintaining the efficiency of MLLMs. Through our contributions, we strive to empower the community to engineer more cost-effective, efficient, and high-performance MLLMs for broad, real-world application scenarios.
@@ 2 Related Work
Advanced language model. The evolution of language models has been marked by significant milestones, starting with early successes like GPT2 radford2019language and BERT devlin2018bert in natural language processing (NLP). These foundational models set the stage for the subsequent development of vastly larger language models, encompassing hundreds of billions of parameters. This dramatic increase in scale has led to the emergence of advanced capabilities as seen in models like GPT-3 brown2020language , Chinchilla hoffmann2022training , OPT zhang2022opt , and BLOOM workshop2022bloom .These large language models (LLMs) have been instrumental in further advancements in the field. For instance, ChatGPT chatgpt and InstructGPT ouyang2022training leverage these powerful models to answer diverse questions and perform complex tasks such as coding. The introduction of open-source LLMs like LLaMA touvron2023llama has further propelled research in this area, inspiring subsequent developments like Alpaca taori2023stanford , Vicuna chiang2023vicuna . These models fine-tune the LLaMA model with additional high-quality instruction datasets, showcasing the versatility and adaptability of LLM frameworks.Among the most notable recent advancements are Phi li2023textbooks and its successor, Phi-2 javaheripi2023phi . These models have demonstrated exceptional performance, rivaling or even surpassing models up to 25 times larger in scale. This indicates a significant shift in the landscape of language modeling, emphasizing efficiency and effectiveness without necessarily relying on sheer size. Such developments mark a new era in the field of NLP, where smaller, more efficient models can achieve results comparable to their much larger counterparts, opening up new possibilities for application and research.
@@
Multimodal language model. In recent years, the trend of aligning visual input to large language models for vision-language tasks has gained significant attention chen2022visualgpt ; tsimpoukelli2021multimodal ; alayrac2022flamingo ; li2023blip2 ; liu2023visual ; liu2023improvedllava ; zhu2023minigpt ; chen2023minigptv2 . Seminal works like VisualGPT chen2022visualgpt and Frozen tsimpoukelli2021multimodal , which utilized pre-trained language models for image captioning and visual question answering. This approach was further advanced by models such as Flamingo alayrac2022flamingo , which incorporated gated cross-attention mechanisms to align pre-trained vision encoders and language models, training on vast image-text pairs. BLIP-2 li2023blip2 introduced an efficient Q-Former for aligning visual and language modalities. These groundbreaking studies have paved the way for further innovations in the field, leading to the development of models like LLaVA liu2023visual and MiniGPT4 zhu2023minigpt , and their subsequent iterations, LLaVA-v1.5 liu2023improvedllava , MiniGPT-v2 chen2023minigptv2 , ArtGPT-4 yuan2023artgpt4 , instruction GPT-4 wei2023instructiongpt4 and Instruction Mining cao2023instruction . These models have demonstrated advanced multimodal capabilities through instruction tuning, showcasing remarkable generalization abilities. Despite their powerful capabilities for visual-language tasks, these multimodal language models often require substantial computational resources. In contrast, TinyGPT-V represents a paradigm shift, harnessing a cost-effective and powerful small language model to achieve a robust, easily deployable model suitable for a variety of real-world vision-language applications. This approach underscores a move towards more efficient yet equally competent multimodal language modeling.
@@ 3 Method
We commence by proposing our vision-language model, TinyGPT-V, then conduct a discussion on the structure of the model and the organization of tasks, and finally introduce the training process for each stage.
@@ 3.1 Model Architecture
In this subsection, we present the architecture of TinyGPT-V, which consists of a visual coder Linear projection layer and a large language model.
Visual encoder backbone. EVA fang2022eva of ViT serves as the visual foundation model in the TinyGPT-V adaptation as same as MiniGPT-v2. The visual foundation remains inactive throughout entire model training. Our model training operates at a picture resolution of 224x224 for Stage 1, 2 and 3 and 448x448 for Stage 4, and we amplify the positional encoding to scale up with an elevated image resolution.
Linear projection layers. The function of the Linear Projection layer is to embed the visual features extracted by the visual encoder into the language model. As well as making an effort to empower extensive language models to comprehend image-based information. Our employment of the Q-Former layer, sourced from the BLIP-2 li2023blip2 architecture, as the initial linear projection layer is driven by the goal to extract maximum functionality from the pre-trained BLIP system when deployed in visual language models. This approach substantially decreases the volume of parameters that require training stages. We utilize linear projection layers initialized with a Gaussian distribution as the second layer. The goal is to bridge the dimensionality gap between the Q-Former output and the language model’s embedding layer, thus, better aligning the visual tokens with the relevant hidden space of the language model. As illustrated in Figure 2, to accelerate the training process of TinyGPT-V, we initially employ the pre-trained Linear projection from MiniGPT-4 (Vicuna 7B) as the foundational layer. Subsequently, we integrate an additional Linear layer projection to effectively bridge into the corresponding hidden space of the Phi-2 model.
Large lanuguage model backbone. We utilize the Phi-2 javaheripi2023phi model as the backbone for our TinyGPT-V large language model. Phi-2 is a 2.7 billion-parameter language model with excellent reasoning and language comprehension, demonstrating state-of-the-art performance among base language models with fewer than 13 billion parameters. In complex benchmarks, Phi-2 matches or outperforms most models 25 times larger. We rely squarely on Phi-2 linguistic tokens to execute several vision language operations. For vision anchoring assignments that require the creation of spatial locations, we explicitly request the linguistic model to generate textual depictions of bounding boxes to signify their geographical coordinates.
@@
Normalization and LoRA for TinyGPT-V. In Section 4.3, we deduce that training smaller large-scale language models for transfer learning, particularly across different modalities (such as from text to image), presents significant challenges. Our investigations reveal that smaller models are particularly susceptible to NaN or INF values during multimodal data computation. This often results in a computational loss value of NaN, causing the initial batch forward propagation to fail. Additionally, a limited number of trainable parameters in these smaller models can contribute to gradient vanishing throughout the training process. To address these issues, as show in Figure 3 (c), we integrate the LLaMA-2 post-norm and input norm mechanisms, implementing RMS Norm after each Multi Head Attention Layer (MHA) to normalize the data for subsequent layers. We also update all Layer norms in the Phi-2 model to enhance training stability, as illustrated in the equation below.
@@
where $x_{post}$ is the input after MHA, $N$ is the number of elements in the vector, and $\epsilon$ is a small constant added for numerical stability.
Furthermore, Henry et al. henry2020querykey have underscored the vital role of Query-Key Normalization in low-resource learning scenarios. Hence, as show in Figure 3 (d), we have incorporated Query-Key Normalization into the Phi-2 model, as detailed in the following equation.
@@
The structure of the LoRA mechanism hu2021lora is show in Figure 3 (a), which is an efficient fine-tuning method in parallel to the frozen pre-training weights as shown in Figure 3 (c), which does not increase the inference time consuming for large language models and is easier to optimise.
3.2 Multi-task Instruction Template
In order to mitigate the potential ambiguity when training a unified multi-modal model to handle various distinct tasks such as visual question answering, image captioning, referring expression comprehension, generation, and object parsing and grounding, we have used the MiniGPT-v2 tokens of task-specific within a multi-task instruction template. It is originating from the LLaMA-2 touvron2023llama conversation template and includes a general input format consisting of image features, a task identifier token, and an instruction input. It have six distinct task identifiers, each correlated to a specific task. For tasks requiring the model to identify spatial locations of referred objects, It utilize textual formatting of bounding boxes, with coordinates normalized within the range of 0 to 100. Overall, the unique task-specific tokens provided by MiniGPT-v2 facilitate disambiguation among tasks, allowing for more precise and accurate task execution.
@@ 3.3 Training Stages
In this subsection the three-stage training process of TinyGPT-V will be described.
Table 1: The full list of datasets used by TinyGPT-V during training.
@@
Caption Text Captions
@@
Warm-up training for the first training stage. During the initial pretraining stage, TinyGPT-V is taught vision-language understanding using a large library of aligned image-text pairs. The model identifies the output from the introduced projection layers as a soft prompt directing it to create relevant texts and to allow large language models to accept inputs from the image modality. The pretraining process uses a dataset combination of Conceptual Caption, SBU and LAION, involving 20000 training steps covering about 5 million image-text pairs.
Pre-training for the second training stage. Following the initial training phase, the large language model becomes equipped to process image modality inputs. To guarantee more consistent performance as the model transitions into the subsequent training stage, we re-employ the dataset from the first stage, specifically for training the LoRA module.
Human-like learning for the third training stage. We fine-tuned this TinyGPT-V model using a selection of image-text pairings from MiniGPT4 or LLaVA, which included instructions like “###Human: <Img><ImageHere></Img> Take a look at this image and describe what you notice.###Assistant:.”. We used a uniform template inclusive of a randomly chosen prompt that improved the model’s capacity for generating responses that were consistent and sounded more natural.
Multi-task learning in the fourth training stage. The fourth training stage of TinyGPT-V focuses on enhancing its conversation ability as a chatbot by tuning the model with more multi-modal instruction datasets as same as MiniGPT-v2. These datasets, as shown in Table 1, include LLaVA, Flickr30k, a mixing multi-task dataset, and Unnatural Instruction. The LLaVA dataset is utilized for multi-modal instruction tuning with detailed descriptions and complex reasoning examples. The Flickr30k dataset is used to improve grounded image caption generation and object parsing and grounding capabilities. Additionally, a mixing multi-task dataset is created to improve the model’s handling of multiple tasks during multi-round conversations. Finally, to recover the language generation ability, the Unnatural Instruction dataset is added to the third-stage training of TinyGPT-V.

### v2

@@ Abstract
In recent years, multimodal large language models (MLLMs) such as GPT-4V have demonstrated remarkable advancements, excelling in a variety of vision-language tasks. Despite their prowess, the closed-source nature and computational demands of such models limit their accessibility and applicability. This study introduces TinyGPT-V, a novel open-source MLLM, designed for efficient training and inference across various vision-language tasks, including image captioning (IC) and visual question answering (VQA). Leveraging a compact yet powerful architecture, TinyGPT-V integrates the Phi-2 language model with pre-trained vision encoders, utilizing a unique mapping module for visual and linguistic information fusion. With a training regimen optimized for small backbones and employing a diverse dataset amalgam, TinyGPT-V requires significantly lower computational resources—24GB for training and as little as 8GB for inference—without compromising on performance. Our experiments demonstrate that TinyGPT-V, with its language model 2.8 billion parameters, achieves comparable results in VQA and image inference tasks to its larger counterparts while being uniquely suited for deployment on resource-constrained devices through innovative quantization techniques. This work not only paves the way for more accessible and efficient MLLMs but also underscores the potential of smaller, optimized models in bridging the gap between high performance and computational efficiency in real-world applications. Additionally, this paper introduces a new approach to multimodal large language models using smaller backbones. Our code and training weights are available in https://github.com/DLYuanGod/TinyGPT-V.
11footnotetext: Zhaoxu is a visiting student in the LAIR lab at Lehigh University.
@@ 1 Introduction
In recent years, the field of artificial intelligence has seen significant advancements through the development of multimodal large language models (MLLMs), such as GPT-4V, which have shown exceptional performance across a range of vision-language tasks (Yang et al., 2023). Despite GPT-4V’s impressive capabilities, its closed-source nature limits its widespread application and adaptability. In contrast, the open-source landscape for MLLMs is rapidly evolving, presenting models like LLaVA and MiniGPT-4 that excel in image captioning (IC), visual question answering (VQA) often comparable GPT-4V in these areas (Dai et al., 2023; Liu et al., 2023a; b; Zhu et al., 2023). Notably, MiniGPT-v2 (Chen et al., 2023) has demonstrated superior performance in various visual grounding and question-answering tasks. However, its training code remains proprietary, which poses challenges for community-driven advancements and adaptability.
Although the impressive vision-language capabilities demonstrated by some open-source MLLMs, they frequently necessitate significant computational resources for training and inference. For example, training LLaVA-v1.5-13B (Liu et al., 2023a) required 8 $\times$ A100 GPUs, each equipped with 80GB of memory, cumulating in 25.5 hours of continuous training. As shown in Figure 1 (a), the underlying performance of large language models, which are integral to MLLMs, is pivotal. Models such as LLaVA-v1.5-13B and MiniGPT-v2 depend on high-capacity backbones which are Vicuna-13b-v1.5 (Zheng et al., 2023) and LLaMA2-7B-Chat (Touvron et al., 2023), respectively, necessitating a substantial number of parameters to effectively tackle complex tasks including IC and VQA.
Figure 1: (a) is the occupancy ratio of each component in MiniGPT-4, and (b) is the occupancy ratio of each component in TinyGPT-V. We have considerably narrowed down the occupancy ratio of the Language model in MLLMs.
Figure 2: Comparison of TinyGPT-V with other current MLLMs models shows that TinyGPT-V achieves cost-effective, efficient, and high-performing with fewer parameters.
We introduce TinyGPT-V, a novel model designed for efficient training and inference, requiring only 24GB of GPU memory for training and as little as 8GB of GPU or CPU memory for inference. This model makes use of the advanced large language model Phi-2 (Javaheripi et al., 2023) and incorporates pre-trained vision modules from (Li et al., 2023a) and CLIP (Radford et al., 2021) as its vision encoder, coupled with a mapping module to facilitate the integration of visual information. During training, TinyGPT-V adopts a novel training methodology focused on small pre-trained backbones, unlike any other MLLMs, utilizing the unique mapping module between the visual encoder and the language model as well as novelty normalization methods, while keeping all other components frozen. For its training dataset, TinyGPT-V employs the multi-tasks datasets, including LAION (Schuhmann et al., 2021), Conceptual Captions (Changpinyo et al., 2021; Sharma et al., 2018), SBU (Ordonez et al., 2011), and others (Lin et al., 2015; Schwenk et al., 2022; Hudson & Manning, 2019; Kiela et al., 2020; Lu et al., 2021; Gurari et al., 2018; Mao et al., 2016; Kazemzadeh et al., 2014; Yu et al., 2016).
In our study, we found that TinyGPT-V exhibits similar traits with GPT-4, especially when doing some VQA and image inference. With only 2.8 billion parameters of its language model, TinyGPT-V employs a unique quantization process, like using 8-bit quantization, making it well-suited for local deployment and inference on 8GB mobile devices. This model represents a significant advancement in achieving a balance between exceptional performance and efficiency in MLLMs, as shown in Figure 2. Our work not only aims to enable the community to develop more cost-effective, efficient, and high-performing MLLMs for widespread real-world applications but also introduces a training framework optimized for small pre-trained backbones.
@@ 2 Related Work
Advanced language model. The evolution of language models has been marked by significant milestones, starting with early successes like GPT2 (Radford et al., 2019) and BERT (Devlin et al., 2018) in natural language processing (NLP). These foundational models set the stage for the subsequent development of vastly larger language models, encompassing hundreds of billions of parameters. This dramatic increase in scale has led to the emergence of advanced capabilities as seen in models like GPT-3 (Brown et al., 2020), Chinchilla (Hoffmann et al., 2022), OPT (Zhang et al., 2022), and BLOOM (Workshop et al., 2022).These large language models (LLMs) have been instrumental in further advancements in the field. For instance, ChatGPT (OpenAI, 2022) and InstructGPT (Ouyang et al., 2022) leverage these powerful models to answer diverse questions and perform complex tasks such as coding. The introduction of open-source LLMs like LLaMA (Touvron et al., 2023) has further propelled research in this area, inspiring subsequent developments like Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023). These models fine-tune the LLaMA model with additional high-quality instruction datasets, showcasing the versatility and adaptability of LLM frameworks.Among the most notable recent advancements are Phi (Li et al., 2023b) and its successor, Phi-2 (Javaheripi et al., 2023). These models have demonstrated exceptional performance, rivaling or even surpassing models up to 25 times larger in scale. This indicates a significant shift in the landscape of language modeling, emphasizing efficiency and effectiveness without necessarily relying on sheer size.
@@
Multimodal language model. In recent years, the trend of aligning visual input to large language models for vision-language tasks has gained significant attention (Chen et al., 2022; Tsimpoukelli et al., 2021; Alayrac et al., 2022; Li et al., 2023a; Liu et al., 2023b; a; Zhu et al., 2023; Chen et al., 2023). Seminal works like VisualGPT (Chen et al., 2022) and Frozen (Tsimpoukelli et al., 2021), which utilized pre-trained language models for image captioning and visual question answering. This approach was further advanced by models such as Flamingo (Alayrac et al., 2022), which incorporated gated cross-attention mechanisms to align pre-trained vision encoders and language models, training on vast image-text pairs. BLIP-2 (Li et al., 2023a) introduced an efficient Q-Former for aligning visual and language modalities. These groundbreaking studies have paved the way for further innovations in the field, leading to the development of models like LLaVA (Liu et al., 2023b) and MiniGPT4 (Zhu et al., 2023), and their subsequent iterations, LLaVA-v1.5 (Liu et al., 2023a), MiniGPT-v2 (Chen et al., 2023), ArtGPT-4 (Yuan et al., 2023), instruction GPT-4 (Wei et al., 2023) and Instruction Mining (Cao et al., 2023). These models have demonstrated advanced multimodal capabilities through instruction tuning, showcasing remarkable generalization abilities.
@@ 3 Method
We briefly introduce our vision-language model, TinyGPT-V, followed by an analysis of its structure, culminating in a detailed description of the training process for each stage.
Figure 4: Architecture of TinyGPT-V. The model takes a visual backbone, which remains frozen during all training phases. We concatenate Q-Former layer visual output tokens from ViT backbone and project them into Phi-2 language model space via two linear projection layers.
@@ 3.1 Model Architecture
In this subsection, we present the architecture of TinyGPT-V, which consists of a visual encoder, projection layers, and a large language model, as shown in Figure 4.
Visual encoder backbone. In the TinyGPT-V, it utilizes EVA (Fang et al., 2022) of the ViT serves as the visual foundation model, which remains inactive during the entire training process. Our model operates at an image resolution of 224x224 for Stages 1, 2, and 3, and at 448x448 for Stage 4. The positional encoding is enhanced to accommodate the increased image resolution which is known as the Relative Position Bias (Dufter et al., 2021). It enhances the model’s understanding of the spatial relationships between elements in an image.
Projection layers. The Projection layers embed visual features extracted by the visual encoder into the language model, enhancing the model’s ability to process image-based information. We adopt the Q-Former layers from the BLIP-2 architecture (Li et al., 2023a) as the initial projection layer, aiming to leverage the full potential of the pre-trained BLIP system within visual language models. This strategy significantly reduces the number of parameters needing training. The second and third layers are linear projection layers, designed to bridge the dimensionality gap between the Q-Former output and the language model’s embedding layer, thereby aligning visual tokens more effectively with the language model’s hidden space. As shown in Figure 6, to expedite TinyGPT-V’s training, we initially use a pre-trained Linear Projection from MiniGPT-4 (Vicuna 7B) as the second layer. We then introduce an additional linear projection layer, initialized with a Gaussian distribution, as the third layer to seamlessly integrate into the hidden space of the Phi-2 model.
Large lanuguage model backbone. Our TinyGPT-V large language model is built upon the Phi-2 model (Javaheripi et al., 2023) as its backbone. Phi-2, a 2.7 billion-parameter language model, exhibits exceptional reasoning and language comprehension abilities, achieving state-of-the-art performance among language models with fewer than 13 billion parameters. In complex benchmarks, Phi-2 either matches or exceeds the performance of models up to 25 times its size. We primarily use Phi-2’s linguistic abilities to do various vision-language tasks. Specifically, for vision reasoning tasks that involve spatial location identification, we instruct the linguistic model to generate textual descriptions of what will happen in the next scenario, representing their objects’ coordinates, as shown in Table 3.
@@
Normalization and LoRA for TinyGPT-V. In Section 4.4, we conclude that training smaller-scale large language models for transfer learning, especially across different modalities (e.g., from text to image), poses significant challenges. Our studies indicate that these smaller models are prone to encountering NaN or INF values during multimodal data computations. This issue often leads to a computational loss value of NaN, thereby causing failure in the initial batch forward propagation. Moreover, the limited number of trainable parameters in these models may lead to gradient vanishing during training. To mitigate these problems, as depicted in Figure 5 (c), we incorporate the post-norm and input norm mechanisms from LLaMA-2, applying RMS Norm after each Multi-Head Attention Layer (MHA) to normalize data for downstream processing. In addition, we have to update all layer norms in the Phi-2 model to improve training stability, as detailed in the subsequent equation.
@@
where $x_{post}$ is the input after MHA, $N$ is the dimension of $x_{post}$.
Furthermore, Henry et al. (2020) have underscored the vital role of Query-Key Normalization in low-resource learning scenarios. Hence, as show in Figure 5 (d), we have incorporated Query-Key Normalization into the Phi-2 model, as detailed in the following equation.
@@
The structure of the LoRA mechanism (Hu et al., 2021) is show in Figure 5 (a), which is an efficient fine-tuning method in parallel to the frozen pre-training weights as shown in Figure 5 (c), which does not increase the inference time consuming for large language models and is easier to optimize.
@@ 3.2 Training Stages
In this subsection, the four-stage training process of TinyGPT-V will be described.
@@
Caption Text Captions, COCO Captions
@@
Table 1: The full list of datasets used by TinyGPT-V during training.
Warm-up training for the first training stage. During the initial pretraining stage, TinyGPT-V is taught vision-language understanding using large datasets of aligned image-text pairs. The model identifies the output from the projection layers as a soft prompt directing it to create relevant texts and to allow large language models to accept inputs from the image modality. The pretraining process uses a dataset combination of Conceptual Caption, SBU, and LAION, involving 20000 training steps covering about 5 million image-text pairs.
Pre-training for the second training stage. Following the initial training stage, the large language model becomes equipped to process image modality inputs. To guarantee more consistent performance as the model transitions into the subsequent training stage, we re-employ the dataset from the first stage, specifically for training the LoRA module.
Instruction tuning for the third training stage. We fine-tuned this TinyGPT-V model using a selection of image-text pairings from MiniGPT4 or LLaVA, which included instructions like “###Human: <Img><ImageHere></Img> Take a look at this image and describe what you notice.###Assistant:.”. We used a uniform template inclusive of a randomly chosen prompt that improved the model’s capacity for generating responses that were consistent and sounded more natural.
Multi-task learning in the fourth training stage. The fourth training stage of TinyGPT-V focuses on enhancing its conversation ability as a chatbot by tuning the model with more multi-modal instruction datasets as shown in Table 1, including LLaVA, Flickr30k, a mixing multi-task dataset, and Unnatural Instruction using multi-tasks template as detailed in appendix A. The LLaVA dataset is utilized for multi-modal instruction tuning with detailed descriptions and complex reasoning examples. The Flickr30k dataset is used to improve grounded image caption generation and object parsing and grounding capabilities. Additionally, a mixing multi-task dataset is created to improve the model’s handling of multiple tasks during multi-round conversations. Finally, to recover the language generation ability, the Unnatural Instruction dataset is added to the third-stage training of TinyGPT-V.